x= 0.12345........998999  digits 1 to 999 used then what is 2014 th digit right to the decimal

Grade:12th pass

1 Answers

Dheeraj
5 years ago
Since no of one digit no  after decimal =9. (123456789)
Now no. Of two digit no are =90
 So no. Of digit =2*90=180
Total(1+2)no of digit=189
 
Now for three digit no 
  Total no of 3 digit no =900
 No of digit =3*900=2700
 
Now we have to find 2014 th digit
  Our required digit lies in 3 digit no because total one and two digit no. have189 digit
Now 2014-189=1825
Is the req. Digit
( 1825-1)/3=608
Now usin ap
608 th term 
=100+607=707
 Next no . =708
So 2014 th digit after decimal will be 7